# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ 7 Conclusions and perspectives
Dimension reduction is becoming increasingly important. PCA is probably the most utilized method in practice due to its simplicity, despite lacking guarantees for its effectiveness as a preprocessing tool for clustering or outlier detection. In these contexts, PP appears much more natural and has theoretical justification Radojičić et al. (2021). However, PP is computationally expensive. From that perspective, ICS is a promising alternative – it is computationally less demanding, and theoretical properties can be derived in quite general mixture model frameworks. As shown in the seminal paper Tyler et al. (2009), ICS can recover the FDS. It essentially reduces to an eigenvalue problem where the noise space has identical eigenvalues. The crucial question is then whether all eigenvalues belonging to the space spanning the FDS are distinct from the noise value. In the two-component Gaussian mixture model using the combination $\mathrm{COV}-\mathrm{COV}_{4}$, it was known that this generally works except for one specific mixing proportion. In this work, we extended this result to cover richer mixture models and different scatter combinations, though theoretical studies seem mainly feasible only for $\mathrm{COV}-\mathrm{COV}_{4}$. Based on the findings, it seems that ICS is indeed a natural dimension reduction method for clustering and outlier detection, and with an increasing number of clusters ($k$ still smaller than $p$), the FDS will be estimated.
@@
In practice, it seems customary to compare the performance of different scatter combinations and choose the best one, which is, however, still done heuristically, and corresponding tools for a comparison could be developed. One crucial issue in practice is then also to establish what the noise space eigenvalue is, and which eigenvalues are distinct from that value. Some heuristic rules are, for example, discussed in Archimbaud et al. (2018); Alfons et al. (2024); Radojicic and Nordhausen (2020), but inferential tools are still missing. So far, only Kankainen et al. (2007) propose some tests using $\mathrm{COV}-\mathrm{COV}_{4}$ if all eigenvalues are equal in the Gaussian case (i.e., testing for multivariate normality) and Luo and Li (2016); Radojicic and Nordhausen (2020); Nordhausen et al. (2022) in a non-Gaussian component analysis framework for the equality of eigenvalues for components belonging to Gaussian components. Similar tests might be of interest also in model (1).
@@
All computations in Sections 2, 3, 4 are performed with Python version 3.10.13 (Van Rossum and Drake, 2009), notable packages include NumPy (Harris et al., 2020) for numerical computations, Pandas (The pandas development team, 2020; McKinney, 2010) for data manipulation and Plotly (Plotly Technologies Inc., 2015) for data visualization in Subsection 2.4, Sections 3 and 4. Furthermore, all simulations in Section 6 are performed with R version 4.3.3 (R Core Team, 2023) and uses the R packages ICS (Nordhausen et al., 2008, 2023) for ICS, ICSClust (Archimbaud et al., 2023b) for selecting the components, rrcov (Todorov and Filzmoser, 2009) for the MCD scatter matrix. Replication files for the theoretical computations and simulations are available upon request.
@@ Acknowledgments
We thank Camille Mondon for stimulating discussions about ICS. Anne Ruiz-Gazen acknowledges funding from the French National Research Agency (ANR) under the Investments for the Future (Investissements d’Avenir) program, grant ANR-17-EURE-0010. Klaus Nordhausen was supported by the HiTEc COST Action (CA21163).
@@ Appendix A.   Calculation details for the scatter pair $\mathrm{COV}-\mathrm{COV}_{4}$
Because of the invariance property of ICS, we consider wlog (see A.2 in Tyler et al., 2009) the mixture :
@@
$X\sim\sum_{j=1}^{k}\alpha_{j}{\cal N}_{p}(t_{j},I_{p})$, where the $t_{j}=(t_{ji})\in\mathbb{R}^{p}$ are distinct for $j\in\{1,\ldots,k\}$ and can be written $\sum_{i=1}^{j}t_{ji}e_{i}$ (with $e_{i}$ the $p$-dimensional vector with one in the $i$-th coordinate and zero elsewhere for $j\in\{1,\ldots,k-1\}$), and $t_{k}$ is the $p$-dimensional zero vector. Furthermore, to ease ourcomputations, we define $X^{\mbox{c}}:=X-\mathbb{E}(X)$ whose distribution is $\sum_{j=1}^{k}\alpha_{j}{\cal N}_{p}(t_{j}^{\mbox{c}},I_{p})$, where $t_{j}^{\mbox{c}}=t_{j}-\mathbb{E}(X)$ has coordinates $t_{ji}^{\mbox{c}}$, for $j\in\{1,\ldots,k\}$ and $i\in\{1,\ldots,q\}$. The covariance matrix of $X$ is $\mathrm{COV}=\Gamma_{W}+\Gamma_{B}$ where $\Gamma_{W}=I_{p}$ is the within-group covariance matrix and $\Gamma_{B}=\sum_{j=1}^{k}\alpha_{j}t_{j}^{\mbox{c}}(t_{j}^{\mbox{c}})^{\top}$ is the between-group covariance matrix. This yields:
$\mathrm{COV}=\begin{bmatrix}\beta&0\\ 0&I_{p-q}\end{bmatrix}\quad\text{and}\quad\mathrm{COV}^{-1}=\begin{bmatrix}B&0\\ 0&I_{p-q}\end{bmatrix},$
@@
where the terms of $\beta$ are $\beta_{ms}=\sum_{j=1}^{k}\alpha_{j}t_{jm}^{\mbox{c}}t_{js}^{\mbox{c}}$ for distinct $m,s\in\{1,\ldots,q\}$, and $\beta_{mm}=1+\sum_{j=1}^{k}\alpha_{j}(t_{jm}^{\mbox{c}})^{2}$ for $m\in~{}\{1,\ldots,q\}$. $B=(b_{ij})$ is $q\times q$ matrix whose terms are difficult to express. Concerning $\mathrm{COV}_{4}$, we get:
$\mathrm{COV}_{4}$ $=\frac{1}{p+2}\times\mathbb{E}\big{(}(X^{\mbox{c}})^{\top}\mathrm{COV}^{-1}X^{\mbox{c}}X^{\mbox{c}}(X^{\mbox{c}})^{\top}\big{)}=\frac{1}{p+2}\times\mathbb{E}\Bigg{[}\Bigg{(}\sum_{i=1}^{q}\sum_{j=1}^{q}x^{\mbox{c}}_{i}x^{\mbox{c}}_{j}b_{ij}+\sum_{i=q+1}^{p}(x^{\mbox{c}}_{i})^{2}\Bigg{)}X^{\mbox{c}}(X^{\mbox{c}})^{\top}\Bigg{]},$
@@
where $x^{\mbox{c}}_{i}$ are the coordinates of $X^{\mbox{c}}$, for $i\in\{1,\ldots,p\}$. We proceed to compute the moments of the coordinates of $X^{\mbox{c}}$ to obtain the final form of $\mathrm{COV}_{4}$. Let $f_{X}(\mu)=\mu$, $f_{X^{2}}(\mu,\sigma)=\mu^{2}+\sigma^{2}$, $f_{X^{3}}(\mu,\sigma)=\mu^{3}+3\mu\sigma^{2}$, and $f_{X^{4}}(\mu,\sigma)=\mu^{4}+6\mu^{2}\sigma^{2}+3\sigma^{4}$ be the functions that give the moments of order 1 to 4 of the univariate normal distribution with mean $\mu$ and variance $\sigma^{2}$. The moments of the mixture are equal to the linear combination of the moments of its components. For $i,j\in\{1,\ldots,q\}$, $x^{\mbox{c}}_{i}$ and $x^{\mbox{c}}_{j}$ are not independent, but for each mixture component, the coordinates are independent and we get the following expressions:
$\mathbb{E}[x^{\mbox{c}^{2}}_{a}]&=\sum_{j=1}^{k}\alpha_{j}f_{X^{2}}(t^{\mbox{c}}_{ja},1),\quad\mathbb{E}[x^{\mbox{c}}_{a}x^{\mbox{c}}_{b}]=\sum_{j=1}^{k}\alpha_{j}f_{X}(t^{\mbox{c}}_{ja})f_{X}(t^{\mbox{c}}_{jb}),\quad\mathbb{E}[x^{\mbox{c}^{4}}_{a}]=\sum_{j=1}^{k}\alpha_{j}f_{X^{4}}(t^{\mbox{c}}_{ja},1),\\ \mathbb{E}[x^{\mbox{c}^{3}}_{a}x^{\mbox{c}}_{b}]&=\sum_{j=1}^{k}\alpha_{j}f_{X^{3}}(t^{\mbox{c}}_{ja},1)f_{X}(t^{\mbox{c}}_{jb}),\quad\mathbb{E}[x^{\mbox{c}^{2}}_{a}x^{\mbox{c}^{2}}_{b}]=\sum_{j=1}^{k}\alpha_{j}f_{X^{2}}(t^{\mbox{c}}_{ja},1)f_{X^{2}}(t^{\mbox{c}}_{jb},1),\\ \mathbb{E}[x^{\mbox{c}^{2}}_{a}x^{\mbox{c}}_{b}x^{\mbox{c}}_{c}]&=\sum_{j=1}^{k}\alpha_{j}f_{X^{2}}(t^{\mbox{c}}_{ja},1)f_{X}(t^{\mbox{c}}_{jb})f_{X}(t^{\mbox{c}}_{jc})\\ \mathbb{E}[x^{\mbox{c}}_{a}x^{\mbox{c}}_{b}x^{\mbox{c}}_{c}x^{\mbox{c}}_{d}]&=\sum_{j=1}^{k}\alpha_{j}f_{X}(t^{\mbox{c}}_{ja})f_{X}(t^{\mbox{c}}_{jb})f_{X}(t^{\mbox{c}}_{jc})f_{X}(t^{\mbox{c}}_{jd}),\end{aligned},$
@@
for distinct $a,b,c,d\in\{1,\ldots,q\}$. For $i>q$ and $j\in\{1,\ldots,p\}$, $x^{\mbox{c}}_{i}$ and $x^{\mbox{c}}_{j}$ are independent which leads to many elements of $\mathrm{COV}_{4}$ being equal to 00 and to the following expression: $\mathrm{COV}_{4}=\begin{bmatrix}\Psi&0\\ 0&I_{p-q}\end{bmatrix}$ where $\Psi=(\psi_{ms})$ is $q\times q$ matrix whose terms for $m,s\in\{1,\ldots,q\}$ are given by:
@@ Appendix A.2.   The case of Dirac mixture in Section 3
Upon the removal of the noise and the dimensions not associated with FDS, the Gaussian distributions are simplified to Dirac distributions with parameter $t_{j}$, denoted by $\delta_{t_{j}}$: $X\sim\sum_{j=1}^{k}\alpha_{j}\delta_{t_{j}},$ where $t_{j}=\sum_{i=1}^{j}t_{ji}e_{i}$, and $e_{i}$ is a $q$-dimensional vector with one in the $i$-th coordinate and zero elsewhere for $j\in\{1,\ldots,k-1\}$, $t_{k}$ is the zero vector, and the $t_{j}$ are distinct for $j\in\{1,\ldots,k\}$.

### v2

@@ 8 Conclusions and perspectives
Dimension reduction is becoming increasingly important. PCA is probably the most utilized method in practice due to its simplicity, despite lacking guarantees for its effectiveness as a preprocessing tool for clustering or outlier detection. In these contexts, PP appears much more natural and has theoretical justification Radojičić et al. (2021). However, PP is computationally expensive. From that perspective, ICS is a promising alternative – it is computationally less demanding, and theoretical properties can be derived in quite general mixture model frameworks. As shown in the seminal paper Tyler et al. (2009), ICS can recover the FDS. It essentially reduces to an eigenvalue problem where the noise space has identical eigenvalues. The crucial question is then whether all eigenvalues belonging to the space spanning the FDS are distinct from the noise value. In the two-group Gaussian mixture model using the combination $\mathrm{COV}-\mathrm{COV}_{4}$, it was known that this generally works except for one specific mixing proportion. In this work, we extended this result to cover richer mixture models and different scatter combinations, though theoretical studies seem mainly feasible only for $\mathrm{COV}-\mathrm{COV}_{4}$. Based on the findings, it seems that ICS is indeed a natural dimension reduction method for clustering and outlier detection, and with an increasing number of clusters ($k$ still smaller than $p$), the FDS will be estimated.
@@
In practice, it seems customary to compare the performance of different scatter combinations and choose the best one, which is, however, still done heuristically, and corresponding tools for a comparison could be developed. One crucial issue in practice is then also to establish what the noise space eigenvalue is, and which eigenvalues are distinct from that value. Some heuristic rules are, for example, discussed in Archimbaud et al. (2018); Alfons et al. (2024); Radojičić and Nordhausen (2020), but inferential tools are still missing. So far, only Kankainen et al. (2007) propose some tests using $\mathrm{COV}-\mathrm{COV}_{4}$ if all eigenvalues are equal in the Gaussian case (i.e., testing for multivariate normality) and Luo and Li (2016); Radojičić and Nordhausen (2020); Nordhausen et al. (2022) in a non-Gaussian component analysis framework for the equality of eigenvalues for components belonging to Gaussian components. Similar tests might be of interest also in model (1).
@@
All computations in Sections 2, 3, 4 are performed with Python version 3.10.13 (Van Rossum and Drake, 2009), notable packages include NumPy (Harris et al., 2020) for numerical computations, Pandas (The pandas development team, 2020; McKinney, 2010) for data manipulation and Plotly (Plotly Technologies Inc., 2015) for data visualization in Subsection 2.4, Sections 3 and 4. Furthermore, all simulations in Section 6 and the analysis of the application in Section 7 are performed with R version 4.3.3 (R Core Team, 2023) and uses the R packages ICS (Nordhausen et al., 2008, 2023) for ICS, ICSClust (Archimbaud et al., 2023b) for selecting the components, rrcov (Todorov and Filzmoser, 2009) for the MCD scatter matrix. Replication files for the theoretical computations and simulations are available from https://github.com/AuroreAA/ICS_FDS-Replication.
@@ Acknowledgments
We thank the editor, associate editor and referees for their constructive remarks. We thank Camille Mondon for stimulating discussions about ICS. Part of Colombe Becquart’s work was done during an Erasmus+ visit funded by the University of Toulouse at the University of Helsinki. Anne Ruiz-Gazen acknowledges funding from the French National Research Agency (ANR) under the Investments for the Future (Investissements d’Avenir) program, grant ANR-17-EURE-0010. Klaus Nordhausen was supported by the HiTEc COST Action (CA21163) and the Research Council of Finland (363261).
@@ Appendix A.   Calculation details for the scatter pair $\mathrm{COV}-\mathrm{COV}_{4}$
Because of the invariance property of ICS, we consider w.l.o.g. (see A.2 in Tyler et al., 2009) the mixture : $\mathbf{X}\sim\sum_{j=1}^{k}\alpha_{j}{\cal N}_{p}(\mathbf{t}_{j},\mathbf{I}_{p})$, where the $\mathbf{t}_{j}=(t_{ji})\in\mathbb{R}^{p}$ are distinct for $j\in\{1,\ldots,k\}$ and can be written $\sum_{i=1}^{j}t_{ji}\mathbf{e}_{i}$ (with $\mathbf{e}_{i}$ the $p$-dimensional vector with one in the $i$-th coordinate and zero elsewhere for $j\in\{1,\ldots,k-1\}$), and $\mathbf{t}_{k}$ is the $p$-dimensional zero vector. Furthermore, to ease our computations, we define $\mathbf{X}^{\mbox{c}}:=\mathbf{X}-\mathbb{E}(\mathbf{X})$ whose distribution is $\sum_{j=1}^{k}\alpha_{j}{\cal N}_{p}(\mathbf{t}_{j}^{\mbox{c}},\mathbf{I}_{p})$, where $\mathbf{t}_{j}^{\mbox{c}}=\mathbf{t}_{j}-\mathbb{E}(\mathbf{X})$ has coordinates $t_{ji}^{\mbox{c}}$, for $j\in\{1,\ldots,k\}$ and $i\in\{1,\ldots,q\}$. The covariance matrix of $\mathbf{X}$ is $\mathrm{COV}=\boldsymbol{\Gamma}_{W}+\boldsymbol{\Gamma}_{B}$ where $\boldsymbol{\Gamma}_{W}=\mathbf{I}_{p}$ is the within-group covariance matrix and $\boldsymbol{\Gamma}_{B}=\sum_{j=1}^{k}\alpha_{j}\mathbf{t}_{j}^{\mbox{c}}(\mathbf{t}_{j}^{\mbox{c}})^{\top}$ is the between-group covariance matrix. This yields:
$\mathrm{COV}=\begin{bmatrix}\boldsymbol{\beta}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix},\quad\mathrm{COV}^{-1}=\begin{bmatrix}\mathbf{B}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix},$
@@
where the terms of $\boldsymbol{\beta}$ are $\beta_{ms}=\sum_{j=1}^{k}\alpha_{j}t_{jm}^{\mbox{c}}t_{js}^{\mbox{c}}$ for distinct $m,s\in\{1,\ldots,q\}$, and $\beta_{mm}=1+\sum_{j=1}^{k}\alpha_{j}(t_{jm}^{\mbox{c}})^{2}$ for $m\in~{}\{1,\ldots,q\}$. $\mathbf{B}=(b_{ij})$ is $q\times q$ matrix whose terms are difficult to express. Concerning $\mathrm{COV}_{4}$, we get:
$\mathrm{COV}_{4}$ $=\frac{1}{p+2}\times\mathbb{E}\big{(}(\mathbf{X}^{\mbox{c}})^{\top}\mathrm{COV}^{-1}\mathbf{X}^{\mbox{c}}\mathbf{X}^{\mbox{c}}(\mathbf{X}^{\mbox{c}})^{\top}\big{)}=\frac{1}{p+2}\times\mathbb{E}\Bigg{[}\Bigg{(}\sum_{i=1}^{q}\sum_{j=1}^{q}x^{\mbox{c}}_{i}x^{\mbox{c}}_{j}b_{ij}+\sum_{i=q+1}^{p}(x^{\mbox{c}}_{i})^{2}\Bigg{)}\mathbf{X}^{\mbox{c}}(\mathbf{X}^{\mbox{c}})^{\top}\Bigg{]},$
where $x^{\mbox{c}}_{i}$ are the coordinates of $\mathbf{X}^{\mbox{c}}$, for $i\in\{1,\ldots,p\}$. We proceed to compute the moments of the coordinates of $\mathbf{X}^{\mbox{c}}$ to obtain the final form of $\mathrm{COV}_{4}$.
@@
Let $f_{X}(\mu)=\mu$, $f_{X^{2}}(\mu,\sigma)=\mu^{2}+\sigma^{2}$, $f_{X^{3}}(\mu,\sigma)=\mu^{3}+3\mu\sigma^{2}$, and $f_{X^{4}}(\mu,\sigma)=\mu^{4}+6\mu^{2}\sigma^{2}+3\sigma^{4}$ be the functions that give the moments of order 1 to 4 of the univariate normal distribution with mean $\mu$ and variance $\sigma^{2}$. The moments of the mixture are equal to the linear combination of the moments of its components. For $i,j\in\{1,\ldots,q\}$, $x^{\mbox{c}}_{i}$ and $x^{\mbox{c}}_{j}$ are not independent, but for each mixture component, the coordinates are independent and we get the following expressions:
$\mathbb{E}[x^{\mbox{c}^{2}}_{a}]&=\sum_{j=1}^{k}\alpha_{j}f_{X^{2}}(t^{\mbox{c}}_{ja},1),\quad\mathbb{E}[x^{\mbox{c}}_{a}x^{\mbox{c}}_{b}]=\sum_{j=1}^{k}\alpha_{j}f_{X}(t^{\mbox{c}}_{ja})f_{X}(t^{\mbox{c}}_{jb}),\quad\mathbb{E}[x^{\mbox{c}^{4}}_{a}]=\sum_{j=1}^{k}\alpha_{j}f_{X^{4}}(t^{\mbox{c}}_{ja},1),\\ \mathbb{E}[x^{\mbox{c}^{3}}_{a}x^{\mbox{c}}_{b}]&=\sum_{j=1}^{k}\alpha_{j}f_{X^{3}}(t^{\mbox{c}}_{ja},1)f_{X}(t^{\mbox{c}}_{jb}),\quad\mathbb{E}[x^{\mbox{c}^{2}}_{a}x^{\mbox{c}^{2}}_{b}]=\sum_{j=1}^{k}\alpha_{j}f_{X^{2}}(t^{\mbox{c}}_{ja},1)f_{X^{2}}(t^{\mbox{c}}_{jb},1),\\ \mathbb{E}[x^{\mbox{c}^{2}}_{a}x^{\mbox{c}}_{b}x^{\mbox{c}}_{c}]&=\sum_{j=1}^{k}\alpha_{j}f_{X^{2}}(t^{\mbox{c}}_{ja},1)f_{X}(t^{\mbox{c}}_{jb})f_{X}(t^{\mbox{c}}_{jc}),\quad\mathbb{E}[x^{\mbox{c}}_{a}x^{\mbox{c}}_{b}x^{\mbox{c}}_{c}x^{\mbox{c}}_{d}]=\sum_{j=1}^{k}\alpha_{j}f_{X}(t^{\mbox{c}}_{ja})f_{X}(t^{\mbox{c}}_{jb})f_{X}(t^{\mbox{c}}_{jc})f_{X}(t^{\mbox{c}}_{jd}),\end{aligned},$
@@
for distinct $a,b,c,d\in\{1,\ldots,q\}$. For $i>q$ and $j\in\{1,\ldots,p\}$, $x^{\mbox{c}}_{i}$ and $x^{\mbox{c}}_{j}$ are independent which leads to many elements of $\mathrm{COV}_{4}$ being equal to 00 and to the following expression: $\mathrm{COV}_{4}=\begin{bmatrix}\boldsymbol{\Psi}&\mathbf{0}\\ \mathbf{0}&\mathbf{I}_{p-q}\end{bmatrix}$ where $\boldsymbol{\Psi}=(\psi_{ms})$ is $q\times q$ matrix whose terms for $m,s\in\{1,\ldots,q\}$ are given by: